Given the equation -4 radical X-3=12 solve for x and identity of it is extraneous.
Answer by stanbon(75887)

Given the equation -4 radical X-3=12 solve for x and identity of it is extraneous.
-------
-4sqrt(x-3) = 12
sqrt(x-3) = -3
-------
No solution by definition.
sqrt(x) is defined as not negative.
Cheers,
Stan H.
---------